OpenStudy (ellie202000):

I really need help. Write the inverse function for the function, f(x) = 1/2x + 4. Then, find the value of f-1(4).

OpenStudy (anonymous):

Okay so to solve for the inverse function, You substitute 'y' for f(x), then switch the y and the x. Then resolve for the y to get the inverse function.

OpenStudy (ellie202000):

y=1/2x+4 then x=1/2y+4 then x-4=1/2y then x-2=y Right?

OpenStudy (anonymous):

uh no

OpenStudy (anonymous):

\[x-4=\frac{1}{2}y\\ 2x-8=y\]

OpenStudy (ellie202000):

Oh I get it thank you so much.

OpenStudy (anonymous):

yw